### v1

@@ 4.2 Neural network architecture
The PINN architecture consists of a feedforward neural network with two hidden layers of 128 neurons each, utilizing hyperbolic tangent (tanh) activation functions. The input layer accepts eight features: temperature_C, pressure_bar, anode_pressure_bar, thickness_um, current_density_A_cm2, membrane_encoded, compression_um, and pt_interlayer_position(for future use). The output layer produces a single value representing hydrogen concentration in mol%.
@@ 4.3 Physics-informed loss function
The total loss function combines data fidelity and physics consistency:
$L_{total}=(1-\beta)L_{data}+\beta L_{physics}$
Where $L_{data}$ represents mean squared error between predictions and experimental measurements, and $L_{physics}$ enforces physical constraints derived from the Omrani et al. (2021) model[bib31]. The physics loss comprises eleven components:
  1. 1.
[…]
  2. 2.
[…]
  3. 3.
    Henry’s law: Hydrogen saturation concentration $c_{H_{2}},sat=S_{H_{2}}\times p_{H_{2}}$,mem with temperature-dependent solubility.
  4. 4.
    Pressure enhancement: Membrane pressure p_H2,m = $\sqrt{(p_{ca^{2}}+K_{D}\times i)}$ accounting for Darcy flow effects.
  5. 5.
    Temperature model: Membrane temperature $T_{m}=T_{stack}+a\times i+b\times i^{2}$ incorporating ohmic heating.
Each physics component is weighted equally within L_physics, with the overall physics weight $\beta$ optimized through systematic investigation (0.01 $\leq\beta\leq$ 1.0).
@@ 4.4 Training procedure
Model training utilized the Adam optimizer[bib48] with initial learning rate 5.5 × 10-3, $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999, and $\epsilon$ = 10-8. Learning rate scheduling via ReduceLROnPlateau decreased the rate by factor 0.5 when validation loss plateaued for 50 epochs, with minimum learning rate 10-6. Batch size was set to 32 samples, with training proceeding for maximum 2,000 epochs.
@@ 4.5 Uncertainty quantification
The ensemble mean provided point predictions while the ensemble standard deviation quantified epistemic uncertainty. For sensitivity analysis, input perturbations followed Gaussian distributions with standard deviations of 10% of the parameter range.
@@ 4.7 Statistical analysis
All statistical analyses were performed using Python 3.10 with NumPy 1.24, SciPy 1.11, and scikit-learn 1.3. Performance metrics (R², MSE, RMSE, MAE, MAPE) were calculated using standard definitions. Significance testing for model comparisons employed paired t-tests with Bonferroni correction for multiple comparisons. Error distributions were assessed for normality using the Shapiro-Wilk test and Q-Q plots.
@@ 4.8 PINN-Physics fusion for extrapolation
ŷfusion = $\alpha$ × ŷPINN + (1-$\alpha$) × ŷphysics
where ŷPINN represents the PINN model output, ŷphysics denotes the prediction from the Omrani et al.[bib31] physics model, and $\alpha$ is the fusion weight (set to 0.5 for equal contribution). This fusion approach was applied exclusively during inference for extrapolation tests, without modification to the PINN training procedure. The physics model parameters were calibrated using the training data to ensure consistency.
@@ A.1 Supplementary Figures
Figure 5: PINN model performance metrics across membrane types. (a-e) Statistical metrics for 184 experimental data points without augmentation: R2 >> 0.987 for all membranes with five exceeding 0.998; RMSE << 0.04% for standard membranes; MSE << 0.002 for five membrane types; MAE within ±0.03% H2; MAPE << 2.1% except Nafion D2021 (5.78%) due to limited data (n = 21). (f) Parity plot showing excellent agreement between predictions and measurements (overall R2 = 0.9984). Red dashed line indicates perfect prediction; gray bands show ±5% error. Membrane-specific clustering confirms successful capture of material transport characteristics. Metrics from 5-fold cross-validation represent conservative estimates.
@@ A.2 Supplementary Tables
Table 2: Detailed Performance Metrics by Dataset. Performance comparison of PINN model versus physics equation for each experimental dataset after data augmentation.
@@
2021_Omrani 198 99.50 98.80 +0.70 0.0587 2.36
2021_Martin 75 99.94 99.31 +0.63 0.0165 2.03
2020_Bernt 370 99.94 99.81 +0.13 0.0479 3.64
2019_Garbe 171 99.99 99.48 +0.51 0.0078 0.40
2025_Wu 300 99.97 99.24 +0.73 0.0594 1.26
@@
Pure NN No fusion 88.3% 65.2% 43.8% 16.63%
PINN ($\beta$=0.3) No fusion 92.1% 73.0% 54.1% 10.08%
PINN ($\beta$=0.3) 50% fusion 98.4% 92.6% 87.3% 6.03%
@@ A.3 Supplementary Analysis: Comparative Analysis with Recent PINN Applications
H2 Crossover (This work) PEM Electrolyzer First PINN for gas crossover; multi-membrane validation R² = 99.87%, RMSE = 0.0348% 184 points $\rightarrow$ 1,114 augmented This work
@@ A.3.3 Computational Efficiency Comparison
A.4 Supplementary Analysis: Economic Analysis of PEM Electrolyzer Implementation
@@
A.4.1 Economic Sensitivity Analysis
@@
A.5 Supplementary Methods
A.5.1 Data Augmentation Protocol
@@ A.5.2 Hyperparameter Optimization
A.5.3 Uncertainty Quantification Details
@@
A.5.4 Physics Loss Components
Detailed equations for each physics constraint:
  1. 1.
    Faraday’s Law:
    $L_{1}$ $=\left\|\dot{M}_{H_{2}}-\frac{i}{2F}\right\|^{2}$ (1)
    $L_{2}$ $=\left\|\dot{M}_{O_{2}}-\frac{i}{4F}\right\|^{2}$ (2)
  2. 2.
    Diffusion Law:
    $L_{3}=\left\|\dot{M}_{co}-\frac{D_{eff}(c_{sat}-c_{an})}{t_{dif}}\right\|^{2}$ (3)
  3. 3.
    Henry’s Law:
    $L_{4}=\left\|c_{sat}-S_{H_{2}}\times p_{H_{2}}\right\|^{2}$ (4)
  4. 4.
    Temperature Model:
    $L_{5}=\left\|T_{m}-(T_{s}+a\times i+b\times i^{2})\right\|^{2}$ (5)
Total physics loss:
$L_{physics}=\frac{1}{5}\sum_{i=1}^{5}L_{i}$ (6)
A.6 Supplementary Discussion
A.6.1 Comparison with Traditional Approaches

### v2

@@ 4.1 Data collection and preprocessing
Dataset definitions: This study utilizes two distinct datasets: (i) original experimental dataset comprising 184 unique hydrogen crossover measurements compiled from eight peer-reviewed publications[bib32, bib6, bib34, bib4, bib36, bib37, bib38, bib39] spanning 2011–2025, and (ii) augmented dataset of 1,114 data points generated through physics-constrained cubic spline interpolation. All performance metrics reported in the main text represent five-fold cross-validation results on the augmented dataset unless explicitly stated otherwise. Comparative analyses between original and augmented datasets are provided in Table 1 and Supplementary Tables 3-4.
@@ 4.2 Neural network architecture
We employed a feedforward neural network with architecture 8→128→128→1 (input→hidden→hidden→output) with tanh activation functions. This shallow architecture was selected based on recent evidence that two-hidden-layer networks provide sufficient approximation capability while avoiding vanishing gradient issues in physics-constrained optimization[bib60, bib61].
@@ 4.3 Physics-informed loss function
The physics-informed neural network enforces fundamental electrochemical transport laws through a composite loss function that seamlessly integrates experimental data fidelity with physical constraints:
$\mathcal{L}_{\text{total}}=(1-\beta)\mathcal{L}_{\text{data}}+\beta\mathcal{L}_{\text{physics}}$ (1)
where $\beta\in[0,1]$ is the physics weight parameter, systematically optimized to $\beta=0.3$ through cross-validation. The physics loss $\mathcal{L}_{physics}$ enforces six fundamental conservation laws governing hydrogen crossover in PEM electrolyzers:
Electrochemical Stoichiometry
Faraday’s law constrains hydrogen and oxygen production rates:
$\mathcal{L}_{F,H_{2}}=\left\|\dot{M}_{H_{2}}-\frac{i}{2F}\right\|^{2}$ (2)
$\mathcal{L}_{F,O_{2}}=\left\|\dot{M}_{O_{2}}-\frac{i}{4F}\right\|^{2}$ (3)
where $i$ denotes current density (A cm-2) and F=96,485F=96{,}485 C mol-1.
Mass Transport
Fick’s law governs diffusive hydrogen crossover through the membrane:
$\mathcal{L}_{\text{Fick}}=\left\|\dot{M}_{H_{2}}^{co}-D_{\text{eff}}^{H_{2}}\frac{c_{\text{sat}}^{ca}-c_{\text{sat}}^{an}}{t_{\text{dif}}}\right\|^{2}$ (4)
with effective diffusion coefficient $D_{eff}^{H_{2}}=(\epsilon_{m}^{w}/\tau_{m})D_{H_{2},w}$ accounting for membrane microstructure.
Thermodynamic Equilibrium
Henry’s law determines hydrogen solubility at electrode-membrane interfaces:
$\mathcal{L}_{\text{Henry}}^{\text{ca}}=\left\|c_{\text{sat}}^{\text{ca}}-S_{H_{2}}^{\text{ca}}\times P_{\text{ca}}^{\text{eff}}\right\|^{2}$ (5)
$\mathcal{L}_{\text{Henry}}^{\text{an}}=\left\|c_{\text{sat}}^{\text{an}}-S_{H_{2}}^{\text{an}}\times P_{\text{an}}\right\|^{2}$ (6)
where $P_{ca}^{eff}=\sqrt{P_{ca}^{2}+K_{D}\cdot i}$ accounts for convection-enhanced transport through porous media.
Thermal Effects
Membrane temperature incorporates ohmic heating:
$\mathcal{L}_{\text{thermal}}=\left\|T_{m}-\left(T_{\text{stack}}+0.5\cdot i+0.1\cdot i^{2}\right)\right\|^{2}$ (7)
The composite physics loss integrates these constraints with equal weighting:
$\mathcal{L}_{\text{physics}}=\frac{1}{11}\sum_{j=1}^{11}\mathcal{L}_{j}$ (8)
This formulation ensures gradient-balanced training while maintaining physical consistency across operating conditions spanning 0.05–5.0 A cm-2, 1–200 bar, and 25–85°C. Five additional constraints governing water dynamics, concentration bounds, and auxiliary transport phenomena are detailed in A.6.4.
@@ 4.4 Training procedure
Model training utilized the Adam optimizer [bib48] with the following hyperparameters: initial learning rate $\eta=5.5\times 10^{-3}$, exponential decay rates for the first and second moment estimates ($\gamma_{1}=0.9$, $\gamma_{2}=0.999$), and numerical stability constant $\varepsilon=10^{-8}$. To avoid notation conflict with the physics weight parameter $\beta$ used throughout this work, we denote Adam’s momentum parameters using distinct symbols. Learning rate scheduling via ReduceLROnPlateau decreased $\eta$ by a factor of 0.5 when validation loss plateaued for 50 epochs, with minimum learning rate $\eta_{\min}=10^{-6}$. Batch size was set to 32 samples, with training proceeding for maximum 2,000 epochs.
@@ 4.5 Uncertainty quantification
The ensemble mean provided point predictions while the ensemble standard deviation quantified epistemic uncertainty. For sensitivity analysis, input perturbations followed typical operational variations or measurement uncertainties: $\pm 1\,^{\circ}\mathrm{C},\ \pm 1\,\mathrm{bar},\ \pm 0.1\,\mathrm{A/cm}^{2},\ \pm 1\,\mu\mathrm{m}$.
@@ 4.7 Statistical analysis
Statistical analyses were performed using Python 3.10 with NumPy 1.24, SciPy 1.11, and scikit-learn 1.3. Model performance was evaluated using coefficient of determination ($R^{2}$), root-mean-square error (RMSE), mean absolute error (MAE), and mean absolute percentage error (MAPE). For extrapolation tests beyond the training pressure range (1–80 bar), models were evaluated at 120, 160, and 200 bar with $n=24$ total test samples.
Paired $t$-tests assessed significance of performance differences between PINN and conventional neural network models on matched test conditions ($\alpha=0.05$). Effect sizes were quantified using Cohen’s $d$ to evaluate practical significance. Normality assumptions were validated via Shapiro-Wilk and D’Agostino $K^{2}$ tests. Non-parametric bootstrap confidence intervals (10,000 resamples) provided distribution-free validation of parametric results. Post-hoc power analysis was conducted to verify sample size adequacy. Pressure-specific performance metrics and detailed statistical diagnostics are provided in Supplementary Information A.4.
@@ 4.8 PINN-Physics fusion for extrapolation
$\hat{y}_{\text{fusion}}=\alpha\times\hat{y}_{\text{PINN}}+(1-\alpha)\times\hat{y}_{\text{physics}}$ (9)
where $\hat{y}_{\text{PINN}}$ represents the PINN model output, $\hat{y}_{\text{physics}}$ denotes the prediction from the physics model, and $\alpha$ is the fusion weight (set to 0.5 for equal contribution). This fusion approach was applied exclusively during inference for extrapolation tests, without modification to the PINN training procedure. The physics model parameters were calibrated using the training data to ensure consistency.
@@ A.1 Supplementary Figures
A.1.1 Dataset Analysis Figures
Figure 4: Dataset Analysis #1: Comprehensive dataset analysis of hydrogen crossover in PEM water electrolysis under varying operating conditions. a, Relationship between H2 crossover concentration and current density across eight independent studies (n = 184 data points). b, Membrane-specific H2 crossover behavior for six membrane types under varying current densities (0.05–5.00 A cm-2). c, Temperature dependence of mean H2 concentration across the operating range of 25–85°C. Error bars represent standard deviation. d, Distribution of H2 concentration as a function of operating pressure (1–200 bar). Box plots show median, interquartile range, and outliers. e, Comparative analysis of mean H2 crossover concentration across studies, highlighting the predominant contribution from Wu 2025. Data compiled from Grigoriev 2011, Schalenbach 2013, Stahler 2020, Trinke 2017, Martin 2021, Garbe 2019, Wu 2025, and Bernt 2020.
@@ A.1.1 Dataset Analysis Figures
Figure 6: Datraset Analysis #3: Operating pressure and temperature effects on hydrogen crossover across independent studies. a, Study-specific mean H2 concentration at discrete pressure conditions (1–200 bar), revealing pressure-dependent crossover behavior across eight studies. b, Pressure-dependent H2 concentration trends for individual studies, with Wu2025 demonstrating the steepest pressure sensitivity across the full range (1–200 bar). c, Temperature-dependent H2 concentration profiles at four operating temperatures (25°C, 60°C, 80°C, 85°C), showing study-specific thermal responses. d, Statistical summary of H2 concentration by study showing mean values and standard deviations. Sample sizes: Bernt2020 (n = 37, 1.85 $\pm$ 1.78%), Garbe2019 (n = 24, 1.65 $\pm$ 0.96%), Grigoriev 2011 (n = 9, 0.58 $\pm$ 0.12%), Martin2021 (n = 21, 0.94 $\pm$ 0.77%), Schalenbach 2013 (n = 9, 0.95 $\pm$ 0.70%), Stahler 2020 (n = 8, 0.60 $\pm$ 0.38%), Trinke 2017 (n = 28, 1.94 $\pm$ 0.95%), Wu2025 (n = 48, 5.14 $\pm$ 3.55%).
[…]
A.1.2 Membrane Comparison Figures
[…]
Figure 9: Membrane Comparison #2: PINN model performance metrics across membrane types. (a-e) Statistical metrics for 184 experimental data points without augmentation: R2 >> 0.987 for all membranes with five exceeding 0.998; RMSE << 0.04% for standard membranes; MSE << 0.002 for five membrane types; MAE within ±0.03% H2; MAPE << 2.1% except Nafion D2021 (5.78%) due to limited data (n = 21). (f) Parity plot showing excellent agreement between predictions and measurements (overall R2 = 0.9984). Red dashed line indicates perfect prediction; gray bands show ±5% error. Membrane-specific clustering confirms successful capture of material transport characteristics. Metrics from 5-fold cross-validation represent conservative estimates.
@@ A.2 Supplementary Tables
Table 2: Detailed Performance Metrics by Dataset. Performance comparison of PINN model versus physics equation for each experimental dataset after data augmentation.00footnotetext: *This dataset includes [bib32, bib6, bib34, bib4] as subsets.
@@
2021_Omrani[bib31]* 198 99.50 98.80 +0.70 0.0587 2.36
2021_Martin[bib36] 75 99.94 99.31 +0.63 0.0165 2.03
2020_Bernt[bib37] 370 99.94 99.81 +0.13 0.0479 3.64
2019_Garbe[bib38] 171 99.99 99.48 +0.51 0.0078 0.40
2025_Wu[bib39] 300 99.97 99.24 +0.73 0.0594 1.26
@@
Pure NN No fusion 88.3% 65.2% 43.4% 16.63%
PINN ($\beta$=0.3) No fusion 92.1% 73.0% 51.0% 12.16%
PINN ($\beta$=0.3) 50% fusion 98.4% 92.1% 86.4% 6.53%
@@ A.3 Supplementary Analysis: Comparative Analysis with Recent PINN Applications
H2 Crossover (This work) PEM Electrolyzer First PINN for gas crossover; multi-membrane validation R² = 99.84%, RMSE = 0.093% 184 points $\rightarrow$ 1,114 augmented This work
@@ A.3.3 Computational Efficiency Comparison
A.4 Supplementary Analysis: Statistical Analysis of Extrapolation Performance
A.4.1 Extrapolation Test Design and Statistical Validation
To rigorously assess the extrapolation capabilities of the PINN compared to conventional NN, we conducted comprehensive statistical analyses on pressure conditions substantially beyond the training domain.
Training conditions:
  • Pressure: 1.0–80.0 bar
  • Current density: 0.07–1.8 A cm-2
  • Membrane: Nafion 117
  • Temperature: 25–85°C
  • Total training samples: 42
Test conditions:
[…]
  • Current density: 0.07–1.00 A cm-2
  • Membrane: Nafion 117
  • Temperature: 25°C
  • Total test samples: 24
A.4.2 Statistical Analysis Framework
We employed multiple statistical tests to ensure robust validation of performance differences:
Paired t-test Analysis
The paired t-test was selected due to the matched-pairs nature of predictions on identical test conditions.
Without fusion (PINN vs. NN models):
  • Mean absolute error difference: $0.294\%$ (95% CI: [0.335,0.251][-0.335,-0.251])
  • $t$-statistic: 13.41-13.41, $p<0.001$
  • Statistical power: >0.99>0.99 at $\alpha=0.05$
With fusion (PINN+Physics with 0.5 weight averaging vs. NN models):
  • Mean absolute error difference: $0.832\%$ (95% CI: [1.198,0.534][-1.198,-0.534])
  • $t$-statistic: 4.81-4.81, $p<0.001$
Normality Assessment
While individual model errors showed non-normal distributions (Shapiro-Wilk $p<0.001$), the paired differences satisfied normality assumptions:
  • Without fusion: $W=0.983$, $p=0.941$
  • With fusion: $W=0.730$, $p<0.001$ (robust $t$-test confirmed results)
Effect Size Analysis
Cohen’s $d$ provided standardized effect size measurements:
  • Without fusion: d=0.171d=-0.171 (negligible effect)
  • With fusion: d=0.608d=-0.608 (medium effect)
The medium effect size with fusion indicates practically significant improvements beyond statistical significance alone.
Bootstrap Confidence Intervals
Non-parametric bootstrap analysis (10,000 resamples) confirmed parametric test results:
Table 10: Bootstrap confidence intervals for model comparison
Configuration PINN MAE (95% CI) NN MAE (95% CI) Difference (95% CI)
Without fusion 1.14 [0.54, 1.89] 1.44 [0.84, 2.18] 0.29-0.29 [0.33-0.33, 0.25-0.25]
With fusion 0.60 [0.29, 0.98] 1.44 [0.84, 2.18] 0.83-0.83 [1.20-1.20, 0.53-0.53]
A.4.3 Pressure-Dependent Performance
Extrapolation performance degradation showed distinct patterns:
Table 11: Model performance metrics at extrapolation pressures
Pressure (bar) NN $R^{2}$ PINN $R^{2}$ PINN+Physics $R^{2}$
120 0.883 0.921 0.984
160 0.652 0.730 0.921
200 0.434 0.510 0.864
[…]
A.4.4 Statistical Power and Sample Size Justification
Post-hoc power analysis indicated:
  • Achieved power: 0.99 for detecting MAE differences of 0.29%
  • Minimum detectable difference: 0.05% at 80% power
  • Sample size adequacy: $n=24$ sufficient for detecting medium effect sizes ($d>0.5$)
Figure 10: Statistical comparison of PINN and NN model performance. (a) Histogram comparison of absolute prediction errors showing PINN (mean = 1.1415) versus NN (mean = 1.4353). (b) Q-Q plot assessing normality of error differences (Shapiro-Wilk: W = 0.9828, p = 9.41e-01). (c) Distribution of paired error differences (PINN - NN) with mean = -0.2939 ± 0.1051. (d) Box plots illustrating median and interquartile ranges for both models. (e) Bootstrap confidence intervals (95%, n = 10,000 iterations) for mean absolute errors showing overlapping confidence intervals. (f,g) Residual plots versus predicted values for PINN and NN models demonstrating heteroscedastic error patterns. (h) Cumulative distribution functions indicating that 95% of PINN predictions have errors below 4.6348 compared to 4.9144 for NN. Paired t-test revealed a statistically significant difference (t = -13.4124, p = 2.33e-12, Cohen’s d = -0.1711, negligible effect; n = 24).
Figure 11: Statistical comparison of PINN(0.5)+Physics(0.5) fusion vs. NN model performance. (a) Histogram comparison of absolute prediction errors showing PINN fusion (mean = 0.6032) versus NN (mean = 1.4353). (b) Q-Q plot assessing normality of error differences (Shapiro-Wilk: W = 0.7303, p = 2.60e-05). (c) Distribution of paired error differences (PINN fusion - NN) with mean = -0.8317 ± 0.8294. (d) Box plots illustrating median and interquartile ranges for both models. (e) Bootstrap confidence intervals (95%, n = 10,000 iterations) for mean absolute errors showing overlapping confidence intervals. (f,g) Residual plots versus predicted values for PINN fusion and NN models demonstrating heteroscedastic error patterns. (h) Cumulative distribution functions indicating that 95% of PINN fusion predictions have errors below 2.4707 compared to 4.9144 for NN. Paired t-test revealed a statistically significant difference (t = -4.8089, p = 7.50e-05, Cohen’s d = -0.6083, medium effect; n = 24).
A.5 Supplementary Analysis: Economic Analysis of PEM Electrolyzer Implementation
@@
A.5.1 Economic Sensitivity Analysis
@@
A.6 Supplementary Methods
A.6.1 Data Augmentation Protocol
@@ A.6.2 Hyperparameter Optimization
A.6.3 Uncertainty Quantification Details
@@
A.6.4 Complete Physics Constraints
The physics-informed loss function incorporates eleven constraints total, with five auxiliary components complementing the six fundamental laws in the main text:
Water Activity and Content Dynamics
$\mathcal{L}_{\text{water\_activity}}=\left\|a-\frac{P_{H_{2}O}}{P_{\text{sat}}}\right\|^{2}$ (10)
$\mathcal{L}_{\text{water\_content}}=\left\|\lambda_{m}-\left(0.043+17.81a-39.85a^{2}+36.0a^{3}\right)\right\|^{2}$ (11)
where water activity $a$ determines membrane hydration state following Springer’s correlation.
Transport Property Constraints
$\mathcal{L}_{\text{diffusion}}=\left\|D_{\text{eff}}^{H_{2}}-\frac{\epsilon_{m}^{w}}{\tau_{m}}D_{H_{2},w}\right\|^{2}$ (12)
Temperature-Dependent Diffusivity
$\mathcal{L}_{\text{temperature\_dep}}=\left\|D_{H_{2},w}-D_{0}(1+\alpha_{T}(T-T_{\text{ref}}))\right\|^{2}$ (13)
Note: A linearized temperature dependence is employed instead of the full Arrhenius equation to enhance numerical stability during training while maintaining accuracy across the operating temperature range (25–85°C). The linear coefficient $\alpha_{T}=0.01$ K-1 was validated against experimental data.
Crossover Concentration
$\mathcal{L}_{\text{crossover}}=\left\|X_{H_{2}}^{\text{ca}}-\frac{\dot{M}_{H_{2}}^{\text{co}}}{\dot{M}_{O_{2}}}\times 100\%\right\|^{2}$ (14)
Implementation Details
The eleven physics constraints are implemented with equal weighting ($w_{j}=1$) in the composite loss function. The linearized temperature model (13) reduces computational complexity from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n)$ compared to the exponential Arrhenius form while maintaining $R^{2}>0.99$ correlation with experimental diffusivity data in the operational range. This simplification enables stable gradient propagation through the 8-128-128-1 neural architecture without sacrificing prediction accuracy.
Physics Parameter Values
  • Faraday constant: F=96,485F=96{,}485 C mol-1
  • Membrane porosity: $\epsilon_{m}^{w}=0.28$
  • Tortuosity factor: $\tau_{m}=1.5$
  • Darcy coefficient: $K_{d}=1.0\times 10^{-5}$ bar2/(A cm-2)
  • Reference temperature: $T_{\text{ref}}=298$ K
  • Base diffusion coefficient: $D_{0}=1.0\times 10^{-9}$ m2/s
These parameters were fixed based on literature consensus values to reduce model complexity while maintaining physical interpretability.
A.7 Supplementary Discussion
A.7.1 Comparison with Traditional Approaches